OpenStudy (ebonyserrano):

Sam drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took 10 hours. When Sam drove home, there was no traffic and the trip only took 7 hours. If his average rate was 18 miles per hour faster on the trip home, how far away does Sam live from the mountains? Do not do any rounding.

OpenStudy (ebonyserrano):

@YanaSidlinskiy @mhchen

OpenStudy (518nad):

@Will.H

OpenStudy (will.h):

i have to go yana is good

OpenStudy (518nad):

#will_is_better

OpenStudy (will.h):

mate 5 mins ago you told me to flutter off make up your mind

OpenStudy (518nad):

its a JOKE

OpenStudy (518nad):

you're the worst

OpenStudy (will.h):

that's better

OpenStudy (yanasidlinskiy):

Yana is Not good😒😒😒😒😒😭😭😭 That was rude, 518

OpenStudy (518nad):

-.- ok ur not that bad... ur just too young

OpenStudy (518nad):

too innocent to see the world for what it is

OpenStudy (will.h):

Yana gurl u r great

OpenStudy (yanasidlinskiy):

Will nahh.. Take over(;

OpenStudy (will.h):

and just ignore 518 he's silly

OpenStudy (ebonyserrano):

he is

OpenStudy (will.h):

Lol

OpenStudy (yanasidlinskiy):

distance = rate * time distance = slowrate * 10 hours distance = fastrate * 7 hours He's rude too. And you can't ignore Will, Ebony. He's sucha smarty. Can't let him go..

OpenStudy (ebonyserrano):

ok :) thanks for helping me

OpenStudy (ebonyserrano):

so would i multiply 10 and 7 and divide?

OpenStudy (yanasidlinskiy):

slowrate * 10 hours = distance = fastrate * 7 hours slowrate * 10 hours = fastrate * 7 hours slowrate * 10 hours = (slowrate + 18 mph) * 7 hours

OpenStudy (ebonyserrano):

x*10=(x+18)*7?

OpenStudy (will.h):

No divide slow rate * 10 by 7 hrs Slow rate 10/7 = Slow rate + 18 mph Good work yana carry on

OpenStudy (ebonyserrano):

when i divided 10 and 7 i got a really long decimal

OpenStudy (yanasidlinskiy):

What is it in a fraction.

OpenStudy (yanasidlinskiy):

There's 2 fractions that we need..

OpenStudy (yanasidlinskiy):

Whatever the 2 fractions are, let me know. I'll get you through it.. I'll use whatever you give me as your answer.

OpenStudy (ebonyserrano):

is it 0.7 or 1.42857

OpenStudy (ebonyserrano):

that was i got when i divided 7 and 10 and 10 and 7

OpenStudy (yanasidlinskiy):

Do you know what those are in fractions?

OpenStudy (ebonyserrano):

7/10

OpenStudy (yanasidlinskiy):

7/10 slowrate = slowrate + 18 mph 10/7 slowrate = 18 mph Slowrate =??

OpenStudy (ebonyserrano):

how would I find the slowrate?

OpenStudy (ebonyserrano):

???

OpenStudy (518nad):

=3

OpenStudy (518nad):

:>

OpenStudy (ebonyserrano):

@Will.H ?

OpenStudy (518nad):

baha ha ha ha ha

OpenStudy (ebonyserrano):

yana are you going to help?

OpenStudy (518nad):

beheheehehe

OpenStudy (518nad):

do u want me help =.=

OpenStudy (ebonyserrano):

no thanks, your explanation will go on for house but thanks for the suggestion

OpenStudy (ebonyserrano):

*hours

OpenStudy (518nad):

Oh my safkl;sdkf; i hate u

OpenStudy (ebonyserrano):

join the club

OpenStudy (518nad):

ill give u an answer like they do

OpenStudy (518nad):

u hate urself?

OpenStudy (ebonyserrano):

i want an explanation though

OpenStudy (518nad):

they dont explain anything.. they just write some silly comments

OpenStudy (ebonyserrano):

you think i just want the answer and i don't i want an explanation

OpenStudy (ebonyserrano):

@Will.H can you help? Yana messaged me and said she's lagging bad

OpenStudy (518nad):

those silly comments dont mean anything lol its the same as anyone would get from looking at solution

OpenStudy (ebonyserrano):

@Will.H ?????

OpenStudy (will.h):

@Ebonyserrano i will be on in 10 mins in the meanwhile @jhonyy9 can help

OpenStudy (ebonyserrano):

okay thanks

OpenStudy (518nad):

D is Distance of trip V is speed of strip in first half D/V = 10 D/(V+18) =7 Solve for D and V D=420, V = 42

OpenStudy (ebonyserrano):

how would i do that? by multiplying?

OpenStudy (518nad):

im sorry the process of teaching someone how to do algebra is tooo long of an explanation

OpenStudy (ebonyserrano):

ok bye

OpenStudy (518nad):

some people dont like that

OpenStudy (ebonyserrano):

k bye

OpenStudy (ebonyserrano):

thanks for trying

OpenStudy (518nad):

sure thing

OpenStudy (ebonyserrano):

@brooke..help00 can you help?

OpenStudy (ebonyserrano):

@jhonyy9

OpenStudy (brooke..help00):

I'm trying- in the mean time let me tag some really good friends of mine who can also help:) @mathmale @luigi0210 @mikey

OpenStudy (ebonyserrano):

thank you :)

OpenStudy (brooke..help00):

Ok so if there was no traffic wouldn't that mean just him? Of so it only took him 7 hours and with traffic 10. His average rate was 18 mph so wouldn't you multiply 18 times 7 and how many miles that was be your answer?

OpenStudy (brooke..help00):

Ok hold on..

OpenStudy (brooke..help00):

@518 don't say anything your really annoying and getting on my nerces

OpenStudy (brooke..help00):

Nerves*

OpenStudy (518nad):

can you please go away brooke

OpenStudy (518nad):

i could care less

OpenStudy (brooke..help00):

Nad stop! Why did you get off your other account anyways?

OpenStudy (518nad):

D is Distance of trip V is speed of strip in first half D/V = 10 D/(V+18) =7 Solve for D and V D=420, V = 42

OpenStudy (will.h):

1) the speed Sam drove when there is traffic 2) the distance drove each way. Let s = the speed in which Sam drove in traffic and s+18 = the speed Sam drove when there is no traffic The distance traveled is simply speed times time. Hence the equation can be set up as: 10s=7(s+18) simplify the right side7s+126 10s = 7s + 126 -7s -7s 3s = 126 Divide both sides by 3. s=42 Therefore, to find the distance in which Sam lives from the mountains is to multiply his/her speed s and the time he/she takes to get there. D=42*10=420 miles. hope that helps

OpenStudy (ebonyserrano):

thank you all except for 518

OpenStudy (will.h):

LOOOL

OpenStudy (518nad):

O M G

OpenStudy (518nad):

o mmmmmmmmmmmmmmmmm g

OpenStudy (518nad):

so annoying

OpenStudy (will.h):

so emotional mate 518 and @Ebonyserrano let me know if you have other question i'll be here for another 5 mins

OpenStudy (518nad):

im not emotional, you're all emos

OpenStudy (518nad):

no you, its just u!! you are the emo

OpenStudy (will.h):

i could care less about others feelings towards me

OpenStudy (518nad):

that is the emoest of emo things to say

OpenStudy (mhchen):

I wonder who's better

OpenStudy (will.h):

@mhchen about what

OpenStudy (518nad):

i didnt start this, will did he came in like oh u arent doing a good job!

OpenStudy (518nad):

when hes a big noob

OpenStudy (will.h):

Lol 518 mate @Ebony tagged me and you told me to go away when i came in

OpenStudy (will.h):

Besides i was right uhm

OpenStudy (518nad):

uhm uhmmm uhhmmm no

OpenStudy (brooke..help00):

518 am I being mean? XD Okay so if you want to msg me stuff say it ok here. Those are some emotions you are going through in the msgs

OpenStudy (will.h):

Lmao brooke

OpenStudy (brooke..help00):

What do you have to hide?

OpenStudy (518nad):

im not hiding anything!